A wire 20 inches long will be cut into two parts, one to be bent to form a square and the other to form a circle. What is the side of the square and the diameter of the circle so that the combined area is least?


1
Expert's answer
2020-12-17T07:27:39-0500

Solution:

l = length of wire


cut

a = length for square

b = length for circle


l = a + b = 20in


side of square = "\dfrac{a}{4}"


area = "( \\dfrac{a}{4}\t)^2= \\dfrac{a^2}{16}"


perimeter of circle : b = 2"\pi"r

r = "\dfrac{b}{2\pi}"


area = "\pi r^2=\pi ( \dfrac{b}{2\pi})^2= \dfrac{b^2}{4 \pi}"


total area ="\dfrac{a^2}{16}+ \dfrac{b^2}{4\pi}"


a = 20 - b


"\\dfrac{(20-b)^2}{16}\t+ \\dfrac{b^2}{4\\pi}"


"\dfrac{400-40b+b^2}{16}+ \dfrac{16+b^2}{4\pi}"


"25-2.5b+ (\\dfrac{1}{16}+ \\dfrac{1}{4\\pi}\t)b^2"


derive:


"-\\dfrac{10}{4}\t+( \\dfrac{1}{8}+ \\dfrac{1}{2\\pi}\t\t)b"


set to 0 and solve:


"-2.5+( \\dfrac{1}{8}+ \\dfrac{1}{2\\pi}\t\t)b=0"


"( \\dfrac{1}{8}\t+ \\dfrac{1}{2\\pi}\t)b= \\dfrac{10}{4}"


"b= \\dfrac{( \\tfrac{10}{4}\t)}{( \\tfrac{1}{8}+ \\tfrac{1}{2\\pi}\t\t)}"


"b= \\dfrac{1}{( \\tfrac{1}{20}\t+ \\tfrac{1}{5\\pi}\t)}"


b = 8.798

a = 20 - 8.798 = 11.202

a = 11.2

b = 8.8

side of square = "\tfrac{a}{4}=2.8"

diameter of circle = 2r = "2( \tfrac{b}{2\pi})=2.8"